• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Lindgren,Bosted,Gilman

    Log entry time 00:49:05 on October 19,1999

    Entry number 28503

    keyword=shift summary

    21:12 Continuing to take data

    22:05 Stopping run to enter Hall A

    22:30 Carbon doors of FPP have been opened for straight through tests.

    22:45 Electron arm moved to 24.184 degrees, will move to 15 for "NL"
    measurements if ep successfull. Otherwise, this is first angle for 3.115
    GeV gamma+p.

    23 :25 Waiting for beam to come back on

    23:35 Beam shows up, taking 1.5 kHz at 3 microA in e d --> p X.

    23:36 e-arm Q3 down and then reset, came back up okay

    Stopped run at midnight. Back into Hall A to remove carbon blocks.